OpenStudy (twizttiez):

What is the solution to the system of equations represented by the two equations? y=3/2x y=-1/2x+4 (???,???)

OpenStudy (twizttiez):

@shinebrightlikeadimon

OpenStudy (anonymous):

give me a sec

OpenStudy (twizttiez):

Ok same i gotta help some one

OpenStudy (twizttiez):

I'm back ^_^

OpenStudy (twizttiez):

Not that it matters xD

OpenStudy (twizttiez):

@mathmate I'm kinda lost could you help?

OpenStudy (mathmate):

I am equally lost. Not sure what problem you are trying to solve!

OpenStudy (twizttiez):

Me either! xD

OpenStudy (twizttiez):

One moment brb i was tagged

OpenStudy (mathmate):

y=3/2x y=-1/2x+4 could mean a linear system of equations: \(y=\frac{3}{2}x\) \(y=-\frac{1}{2}x+4\) or a different system of linear equations: \(y=\frac{3}{2x}\) \(y=-\frac{1}{2x}+4\)

OpenStudy (twizttiez):

Ok nooow that i know that lol what can i do to solve this?

OpenStudy (mathmate):

or even \(y=-\frac{1}{2x}+4\) \(y=-\frac{1}{2x+4}\)

OpenStudy (twizttiez):

One sec brb

OpenStudy (mathmate):

Your lack of parentheses makes your question ambiguous. All fractions have implicit parentheses for each of numerator and denominator.

OpenStudy (twizttiez):

Thats what the ? is one second let me look at it again

OpenStudy (twizttiez):

@mathmate

OpenStudy (twizttiez):

@Fransisco_w

OpenStudy (twizttiez):

I tagged him to see if he knew what to do

OpenStudy (mathmate):

It would not be ambiguous if you wrote y=(3/2)x y=-(1/2)x+4 Whenever you have y= and y= for both equations, just equate the right hand sides and solve for x. In this case, (3/2)x=-(1/2)x+4 and solve for x. After that, substitute the value of x into either equation and solve for y.

OpenStudy (twizttiez):

So then would it be (3,2)

OpenStudy (twizttiez):

Or no I think it would be (2,3)

OpenStudy (mathmate):

Ok, yes. I thought you put them backwards! Good, well done!

OpenStudy (mathmate):

(2,3) is right.

OpenStudy (twizttiez):

Yay!

OpenStudy (mathmate):

:)